OpenStudy (anonymous):

please answer the 16th question in the question paper below

OpenStudy (anonymous):

pls....answer the 16th question in the paper

OpenStudy (anonymous):

Just put the value of the x in the function \[(5-2\sqrt{6})^2+\frac{1}{(5-2\sqrt{6})^2}\] and simplify.